OpenStudy (arihangdu):

Use a graph calculator to solve the equation -3 cos t = 1 in the interval from 0 less then or equal to theta less then or equal 2 pi. Round to the nearest hundredths. Please help me will madels and fan.

OpenStudy (anonymous):

use the calculator like it say s

OpenStudy (kxttxn):

Here's an online graphing calculator if you don't have one! https://www.desmos.com/calculator

OpenStudy (arihangdu):

I don't know how to do this problem please help me

OpenStudy (arihangdu):

I have it but I don't know how to use.

OpenStudy (arihangdu):

When I drop this equation on calculator it says =0

OpenStudy (arihangdu):

I'm really stuck with this please help me.

OpenStudy (arihangdu):

Please please please help me

OpenStudy (anonymous):

look at the link, it is there

OpenStudy (arihangdu):

Where it is ?

jimthompson5910 (jim_thompson5910):

@kxttxn posted a great graphing calculator link https://www.desmos.com/calculator I also recommend you use it. Even if you have a TI graphing calculator, desmos is a much better option. Here's how you would use it First focus on the original equation. Take the `-3cos(t) = 1` equation and subtract 1 from both sides to get `-3cos(t) - 1 = 0` Now replace every 't' with 'x' to get `-3cos(x)-1 = 0` The next step is to type in `-3cos(x)-1` into the desmos calculator (see the image attachment) Once the graph is set up, you are looking for the points where the curve crosses the x axis (the horizontal axis). You can click on each point to have the coordinates pop up. The first coordinate of each x intercept is going to be part of the solution set. There are infinitely many solutions in the solution set. If you restrict t to be on the interval \(\Large 0 \le t < 2\pi\), then there are 2 solutions.

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (arihangdu):

Please help me to understand this.

OpenStudy (arihangdu):

Hi friend do you know this how to do?

OpenStudy (xapproachesinfinity):

did you what @jim_thompson5910 said

OpenStudy (xapproachesinfinity):

first you write the equation in the desmos -3cosx-1

OpenStudy (arihangdu):

Jim said me but I don't understand how to do this.

OpenStudy (arihangdu):

When I dropping equation on Desmon it gives me the graph ,but here says 0 less then or equal to theta less then or equal to 2pi? What is means?

OpenStudy (xapproachesinfinity):

First step: write the equation here as shown here http://prntscr.com/9gldq9 Second step: points that cross the x axis are the solution of the equation -3cosx-1=0 http://prntscr.com/9gleh9 third step: go to settings http://prntscr.com/9glerp fourth step : change <x< http://prntscr.com/9glfq2 last step: click on the graph and you will see two solution http://prntscr.com/9glgbp

OpenStudy (arihangdu):

🎄 happy Christmas to you. Bye thanksbyee